Consider the following structural model

C1 = ao+ a1Y1

I1=bo +b1Y1+b2Y1-4

Y1=C1+I1+G1

Find the reduced form of the model, C and Y


1
Expert's answer
2021-02-19T13:32:53-0500
"Y=\frac {b_0-4+a_0+G_1}{1-b_1-b_2-a_1}"


"C=a_0+a_1\times \frac {b_0-4+a_0+G_1}{1-b_1-b_2-a_1}"